43. To reach the stoichiometric point in the titration of 25.0 mL of a 0.200 M aqueous solution of HCl, the addition of 20.0 mL of a 0.250 M aqueous solution of NaOH is required. If an additional 1.0 mL of the titrant is added, what is the pH of the analyte solution at this point? A. 11.98 B. 2.26 C. 11.74 D. 13.40 E. 10.40
Answer #1

concentration of excess NaOH(Titrant) = 1*0.25/(25+20+1) = 0.00543 M

pH = 14 -(- log(OH-))

    = 14 - (-log(0.00543))

    = 11.74

answer: C